OpenStudy (anonymous):

Find the domain of the function g(x)= x+9/x^2-1

OpenStudy (anonymous):

All real numbers not equal to 1, -1. Since N/0 = infinity or indeterminant, we find the zeroes of the bottom. Which are 1 and -1.

OpenStudy (tkhunny):

Unclear: \(g(x) = x + \dfrac{9}{x^{2}} - 1\) \(x\ne 0\) \(g(x) = \dfrac{x+9}{x^{2}-1}\) \(x\ne 1\;and\;x\ne -1\)

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (tkhunny):

I guess we'll never know he correct answer.